# Magnetisation moment of a bounded 3D sample: asymptotic recovery from planar measurements on a large disk

### v4

@@ 4. Numerical validation and practical considerations
$j$ $1$ $2$ $3$ $4$
$x_{1}^{\left(j\right)}$ $3.5\cdot 10^{-5}$ m 00 m $4.0\cdot 10^{-5}$ m $-4.0\cdot 10^{-5}$ m
$x_{2}^{\left(j\right)}$ $3.0\cdot 10^{-5}$ m 00 m $-5.5\cdot 10^{-5}$ m $5.5\cdot 10^{-5}$ m
@@
$m_{1}^{\left(j\right)}$ $4.5\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$ $2.5\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$ $-3.0\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$ $-1.0\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$
@@
Finally, in Figure 4.6, we directly test the estimates of the net moment components when the magnetic field is contaminated by noise (with the noise model described above). We observe the persistence of the lower-order estimates for the net moment components whereas the higher-order estimates clearly perform significantly worse.
Let us now briefly comment on the fact that we chose to illustrate the results on a magnetisation distribution with a singular support. Besides its simplicity, this choice is physically motivated as any magnetisation can be thought of a combination of dipole sources. From mathematical (numerical) viewpoint, a magnetic field produced by continuous magnetisation distribution is given by the integral whose numerical approximation (quadrature rule) is nothing but a weighted sum of dipoles. Consequently, we do not expect results for continuous magnetisation distributions to be of any drastical difference. On the other hand, this highlights the applicability of our methodology to the magnetisations that could be much more singular than smooth or square-integrable functions.
@@
Figure 4.2. Estimates of the tangential net moment components $m_{1}$ (left) and $m_{2}$ (right) versus $A$.
Figure 4.3. Estimates of the normal net moment component $m_{3}$ versus $A$: using $x_{1}$ (left) and $x_{2}$ (right) formulas.
Figure 4.4. Convergence for the estimates of $m_{1}$ (left) and $m_{2}$ (right).
@@
Figure 4.6. Estimates of the tangential net moment components $m_{1}$ (left) and $m_{2}$ (right) versus $A$. Noisy data.
Figure 4.7. Estimates of the normal net moment component $m_{3}$ versus $A$: using $x_{1}$ (left) and $x_{2}$ (right) formulas. Noisy data.
@@ 5. Discussion and conclusion
In future works, we should study relation of our asymptotic results to stable estimates of the net moment using constrained optimisation approaches, in particular those obtained in [2] for planar $L^{2}$ magnetisations without the asymptotic assumption. The shape of auxiliary functions to be integrated against the measured field to produce the net moment estimates looks similar to ours. For example, for estimates of $m_{1}$ of second order and higher in Theorem 2.1, this auxiliary function is seen to consist of the main trend (the $2x_{1}$ term) and a correction term (a polynomial depending in the "stretched" variable $x_{1}/A$) which is the largest towards the border of the disk. It would be curious to derive an explicit asymptotic from an integro-differential equation of [2] governing the optimal solution in a particular class. Studying other ways of stabilisation of the estimates and regularisation is also important. In this respect, it is natural to ask a question: what is the best way of using the redundancy of the set of asymptotic formulas for the net moment components to arrive at an estimate with the minimal effect of noise or to have the fastest (non-asymptotic) convergence to the true value of a net moment? For instance, in the present approach, we have a natural redundancy for all estimates of $m_{3}$ of order $3$ and higher due to the freedom of choice $x_{j}=x_{1}$ and $x_{j}=x_{2}$ in formulas (2.6), (2.8) and so on. In Figure 4.3, these estimates give almost indistinguishable results (it can be checked that the difference between the two is not zero but very small), and hence may not look directly useful. However, since having more relations than unknown quantities is always better when dealing with noise, this is still an advantage. Also, it is clear from Step 3 of the proof of Theorem 2.1 that, upon involving higher order polynomials, a richer set of lower-order estimates could be obtained.
@@
The form of asymptotic expressions suggests the direct use of polynomial (in a stretched variables $x_{j}/A$, $j\in\left\{1,2\right\}$) ansantz for the derivation of estimates of higher order, with a potential use of deep learning (an idea suggested by an anonymous reviewer).
@@
Another path for a possible future work is to explore the possibility of obtaining asymptotic expansions analogous to those given in Theorem 2.1 but relying on the smallness of a slghtly different asymptotic parameter. Namely, if, instead of $\left(x_{1}^{2}+x_{2}^{2}\right)^{-3/2}$ and $\left(x_{1}^{2}+x_{2}^{2}\right)^{-5/2}$ in (3.21) and (3.22), respectively, one factored out $\left(x_{1}^{2}+x_{2}^{2}+h^{2}\right)^{-3/2}$ and $\left(x_{1}^{2}+x_{2}^{2}+h^{2}\right)^{-5/2}$ and proceeded with appropriate modifications, the final asymptotic results would have a larger area of validity than that described by (2.1). In particular, this would cover a reduction to a dipolar case in a situation when the measurement area $D_{A}$ is not necessarily large but the value of $h$ is. However, since in the mentioned experimental set-up, the height $h$ is small, this modification would yield a little practical benefit but would complicate analysis of Fourier integrals.
@@
Finally, the results of this work naturally connect to the issue of the asymptotic field extension. Indeed, the asymptotic field expansion at infinity (3.24) is seen to feature $m_{3}$ at the leading order and quantities $c_{1,0}$, $c_{0,1}$ at the next order. While the estimates of $m_{3}$ are given, to a different order, in (2.4), (2.6) and (2.8), it is evident from the proof of Theorem 2.1, that the quantities $c_{1,0}$, $c_{0,1}$ can also be asymptotically estimated, see, e.g., (3.85) and (3.87) (and their versions with $x_{1}$ replaced by $x_{2}$, as well as their lower-order analogs). Thus, this furnishes an explicit $3$-term expansion of $B_{3}$ at infinity which, in general, when solving inverse magnetisation problems, should serve as a better alternative to a simple prolongation of the field by zero outside of the measurement area. Such a strategy can also potentially be used in setting up an iterative scheme. Of course, all of this is meaningful only when the actual measurement area is already large enough so that the asymptotic estimates for $m_{3}$, $c_{1,0}$, $c_{0,1}$ are sufficiently accurate.
@@ Acknowledgements
The Bessel function $J_{n}$ of order $n\in\mathbb{Z}$ is an entire function satisfying the differential equation [19, (10.2.1)]
@@
We have the following series expansion [19, (10.2.2)]
@@
which, due to the entire character of $J_{n}$, is absolutely convergent for every $z\in\mathbb{C}$. Here, $\Gamma$ denotes the Euler gamma function, for which we have, in particular, $\Gamma\left(k\right)=\left(k-1\right)!$ for $k\in\mathbb{N}_{+}$. Moreover, it is worth noting that $01/\Gamma\left(k\right)=$ 0 for $\+𝑘\subscriptk\in\mathbb{Z}\backslash\mathbb{N}_{+}$, which implies vanishing of negative powers of $z$ in expansion (A.2) even for negative orders $n$.
The following integral representation holds [19, (10.9.1)]
@@
[…]
@@
For $x\gg 1$, $n\in\mathbb{N}_{0}$, the leading order asymptotics reads [19, (10.17.2)]
@@
where the estimate of the remainder term is due to the discussion in [19, Sect. 10.17(iii)].
The Bessel functions $J_{n}$, $n\in\mathbb{Z},$ satisfy the connection formula [19, (10.4.1)]
@@
as well as simple recurrence relations [19, (10.6.1)]
@@
The Struve function $H_{n}$ of order $n\in\mathbb{N}_{0}\cup\left\{-1\right\}$ is an entire function defined by the absolutely convergent power series [19, (11.2.1)]
@@
The companion Struve function $K_{n}$ of order $n\in\mathbb{Z}$ is defined [19, (11.2.5)] as
@@
For $x\gg 1$, $n\in\mathbb{N}_{0}$, the following asymptotics hold true [19, (11.6.1)]
@@
and the remainder terms are discussed in [19, Sect. 11.6(i)] and [19, Sect. 10.17(iii)], respectively. The asymptotic behavior of $H_{n}\left(x\right)$ for $x\gg 1$, $n\in\mathbb{N}_{0}$, hence follows from (A.12)–(A.14). In particular, for $x\gg 1$,
@@
First, upon integration by parts (using the asymptotic behavior at infinity of $J_{1}$ given by (A.6)), we have
@@
Here, in the second line, we eliminated the intergral term involving $J_{1}^{\prime\prime}$ using (A.20).
@@
Here, in the first line, we employed integration by parts (together with the asymptotic behavior at infinity of $J_{0}$ given by (A.6)). To arrive at the second line, we used the identity
@@
[…]
@@
[…]
@@
[…]
@@
To deduce relation (A.27), we use [19, (10.22.2)] to write
[…]

### v5

@@ 4. Numerical validation and practical considerations
$j$ 11 22 33 44
$x_{1}^{\left(j\right)}$ $3.5\cdot 10^{-5}$ m 0 m $4.0\cdot 10^{-5}$ m $4.0\cdot 10^{-5}$ m
$x_{2}^{\left(j\right)}$ $3.0\cdot 10^{-5}$ m 0 m $5.5\cdot 10^{-5}$ m $5.5\cdot 10^{-5}$ m
@@
$m_{1}^{\left(j\right)}$ $4.5\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$ $2.5\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$ $3.0\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$ $1.0\cdot 10^{-12}$ $\text{A}\cdot\text{m}^{2}$
@@
Finally, in Figures 4.64.7, we directly test the estimates of the net moment components when the magnetic field is contaminated by noise (with the noise model described above). We observe the persistence of the lower-order estimates for the net moment components whereas the higher-order estimates clearly perform significantly worse.
Let us now briefly comment on the fact that we chose to illustrate the results on a magnetisation distribution with a singular support. Besides its simplicity, this choice is physically motivated as any magnetisation can be thought of a combination of dipole sources. From mathematical (numerical) viewpoint, a magnetic field produced by continuous magnetisation distribution is given by the integral whose numerical approximation (quadrature rule) is nothing but a weighted sum of dipoles. Consequently, we do not expect results for continuous magnetisation distributions to be of any drastic difference. On the other hand, this highlights the applicability of our methodology to the magnetisations that could be much more singular than smooth or square-integrable functions.
@@
Figure 4.2. Estimates of the tangential net moment components $m_{1}$ (left) and $m_{2}$ (right) versus AA.
Figure 4.3. Estimates of the normal net moment component $m_{3}$ versus AA: using $x_{1}$ (left) and $x_{2}$ (right) formulas.
Figure $4.$ Convergence for the estimates of $m_{1}$ (left) and $m_{2}$ (right).
@@
Figure 4.6. Estimates of the tangential net moment components $m_{1}$ (left) and $m_{2}$ (right) versus AA. Noisy data.
Figure 4.7. Estimates of the normal net moment component $m_{3}$ versus AA: using $x_{1}$ (left) and $x_{2}$ (right) formulas. Noisy data.
@@ 5. Discussion and conclusion
In future works, we should study relation of our asymptotic results to stable estimates of the net moment using constrained optimisation approaches, in particular those obtained in [2] for planar $L^{2}$ magnetisations without the asymptotic assumption. The shape of auxiliary functions to be integrated against the measured field to produce the net moment estimates looks similar to ours. For example, for estimates of $m_{1}$ of second order and higher in Theorem 2.1, this auxiliary function is seen to consist of the main trend (the $2x_{1}$ term) and a correction term (a polynomial depending in the "stretched" variable $x_{1}/A$) which is the largest towards the boundary of the disk. It would be curious to derive an explicit asymptotic from an integro-differential equation of [2] governing the optimal solution in a particular class. Studying other ways of stabilisation of the estimates and regularisation is also important. In this respect, it is natural to ask a question: what is the best way of using the redundancy of the set of asymptotic formulas for the net moment components to arrive at an estimate with the minimal effect of noise or to have the fastest (non-asymptotic) convergence to the true value of a net moment? For instance, in the present approach, we have a natural redundancy for all estimates of $m_{3}$ of order $3$ and higher due to the freedom of choice $x_{j}=x_{1}$ and $x_{j}=x_{2}$ in formulas (2.6), (2.8) and so on. In Figure 4.3, these estimates give almost indistinguishable results (it can be checked that the difference between the two is not zero but very small), and hence may not look directly useful. However, since having more relations than unknown quantities is always better when dealing with noise, this is still an advantage. Also, it is clear from Step 3 of the proof of Theorem 2.1 that, upon involving higher order polynomials, a richer set of lower-order estimates could be obtained.
@@
The form of asymptotic expressions suggests the direct use of polynomial (in a stretched variables $x_{j}/A$, $j\in\left\{1,2\right\}$) ansatz for the derivation of estimates of higher order, with a potential application of machine learning for identification of their coefficients.
@@
Another path for a possible future work is to explore the possibility of obtaining asymptotic expansions analogous to those given in Theorem 2.1 but relying on the smallness of a slightly different asymptotic parameter. Namely, if, instead of $\left(x_{1}^{2}+x_{2}^{2}\right)^{-3/2}$ and $\left(x_{1}^{2}+x_{2}^{2}\right)^{-5/2}$ in (3.21) and (3.22), respectively, one factored out $\left(x_{1}^{2}+x_{2}^{2}+h^{2}\right)^{-3/2}$ and $\left(x_{1}^{2}+x_{2}^{2}+h^{2}\right)^{-5/2}$ and proceeded with appropriate modifications, the final asymptotic results would have a larger area of validity than that described by (2.1). In particular, this would cover a reduction to a dipolar case in a situation when the measurement area $D_{A}$ is not necessarily large but the value of $h$ is. However, since in the mentioned experimental set-up, the height $h$ is small, this modification would yield a little practical benefit but would complicate analysis of Fourier integrals.
@@
The results of this work naturally connect to the issue of the asymptotic field extension. Indeed, the asymptotic field expansion at infinity (3.24) is seen to feature $m_{3}$ at the leading order and quantities $c_{1,0}$, $c_{0,1}$ at the next order. While the estimates of $m_{3}$ are given, to a different order, in (2.4), (2.6) and (2.8), it is evident from the proof of Theorem 2.1, that the quantities $c_{1,0}$, $c_{0,1}$ can also be asymptotically estimated, see, e.g., (3.85) and (3.87) (and their versions with $x_{1}$ replaced by $x_{2}$, as well as their lower-order analogs). Thus, this furnishes an explicit $3$-term expansion of $B_{3}$ at infinity which, in general, when solving inverse magnetisation problems, should serve as a better alternative to a simple prolongation of the field by zero outside of the measurement area. Such a strategy can also potentially be used in setting up an iterative scheme. Of course, all of this is meaningful only when the actual measurement area is already large enough so that the asymptotic estimates for $m_{3}$, $c_{1,0}$, $c_{0,1}$ are sufficiently accurate.
Finally, on the practical side, it is essential to test the obtained results on a concrete application to see the actual improvement in the existing results and, more importantly, to address the cases which were not amenable for treatment with previous methodologies. Most natural context would certainly be that of estimation of the remanent magnetisation in (terrestrial or extraterrestrial) rock samples using scanning microscopy data, a setting which was the main motivation of the present work. However, due to close proximity of the setting to the active area of magnetic prospection, see, e.g., a very recent work [9], it will be great to test the already derived formulas in that context. But also, since in that case all three components of the field are available instead of only one, it is worthwhile to investigate improvements of the presented results combining integrals of $B_{1}$, $B_{2}$ and $B_{3}$ giving higher-order estimates of the net moment $\vec{m}$.
@@ Acknowledgements
The Bessel function $J_{n}$ of order $n\in\mathbb{Z}$ is an entire function satisfying the differential equation [20, (10.2.1)]
@@
We have the following series expansion [20, (10.2.2)]
@@
which, due to the entire character of $J_{n}$, is absolutely convergent for every $z\in\mathbb{C}$. Here, $\Gamma$ denotes the Euler gamma function, for which we have, in particular, $\Gamma\left(k\right)=\left(k-1\right)!$ for $k\in\mathbb{N}_{+}$. Moreover, it is worth noting that $1/\Gamma\left(k\right)=0$ for $k\in\mathbb{Z}\backslash\mathbb{N}_{+}$, which implies vanishing of negative powers of $z$ in expansion (A.2) even for negative orders $n$.
The following integral representation holds [20, (10.9.1)]
@@
(A.4) $J_{0}\left(x\right)=\dfrac{1}{\pi}\int_{0}^{\pi}\cos\left(x\sin t\right)\mathrm{d}t=\dfrac{1}{\pi}\int_{0}^{\pi}\cos\left(x\cos t\right)\mathrm{d}t=\frac{1}{2\pi}\int_{0}^{2\pi}\cos\left(x\cos t\right)\mathrm{d}t,\hskip 10.00002ptx\in\mathbb{R},$
@@
For $x\gg 1$, $n\in\mathbb{N}_{0}$, the leading order asymptotics reads [20, (10.17.2)]
@@
where the estimate of the remainder term is due to the discussion in [20, Sect. 10.17(iii)].
The Bessel functions $J_{n}$, $n\in\mathbb{Z},$ satisfy the connection formula [20, (10.4.1)]
@@
as well as simple recurrence relations [20, (10.6.1)]
@@
The Struve function $H_{n}$ of order $n\in\mathbb{N}_{0}\cup\left\{-1\right\}$ is an entire function defined by the absolutely convergent power series [20, (11.2.1)]
@@
The companion Struve function $K_{n}$ of order $n\in\mathbb{Z}$ is defined [20, (11.2.5)] as
@@
For $x\gg 1$, $n\in\mathbb{N}_{0}$, the following asymptotics hold true [20, (11.6.1)]
@@
and the remainder terms are discussed in [20, Sect. 11.6(i)] and [20, Sect. 10.17(iii)], respectively. The asymptotic behaviour of $H_{n}\left(x\right)$ for $x\gg 1$, $n\in\mathbb{N}_{0}$, hence follows from (A.12)–(A.14). In particular, for $x\gg 1$,
@@
First, upon integration by parts (using the asymptotic behaviour at infinity of $J_{1}$ given by (A.6)), we have
@@
Here, in the second line, we eliminated the integral term involving $J_{1}^{\prime\prime}$ using (A.20).
@@
Here, in the first line, we employed integration by parts (together with the asymptotic behaviour at infinity of $J_{0}$ given by (A.6)). To arrive at the second line, we used the identity
@@
Expressions (A.24) and (A.25) imply that the integral on the left-hand side of (A.23) is expressible in terms of two integral quantities: $\int_{\rho}^{\infty}J_{0}\left(x\right)dx$ and $\int_{\rho}^{\infty}J_{2}\left(x\right)dx$. Let us now show that these quantities are simply related:
(A.26) $\int_{\rho}^{\infty}J_{2}\left(x\right)dx=\int_{\rho}^{\infty}J_{0}\left(x\right)dx+2J_{1}\left(\rho\right),$
@@
(A.27) $\int_{\rho}^{\infty}J_{0}\left(x\right)dx=1-\rho J_{0}\left(\rho\right)-\frac{\pi}{2}\rho\left[J_{1}\left(\rho\right)H_{0}\left(\rho\right)-J_{0}\left(\rho\right)H_{1}\left(\rho\right)\right].$
@@
$=$ $\int_{\rho}^{R}J_{0}\left(x\right)dx-2\left(J_{1}\left(R\right)-J_{1}\left(\rho\right)\right),$
@@
To deduce relation (A.27), we use [20, (10.22.2)] to write
(A.29) $\int_{\rho}^{R}J_{0}\left(x\right)dx$ $=\frac{\pi}{2}R\left[J_{0}\left(R\right)H_{-1}\left(R\right)-J_{-1}\left(R\right)H_{0}\left(R\right)\right]-\frac{\pi}{2}\rho\left[J_{0}\left(\rho\right)H_{-1}\left(\rho\right)-J_{-1}\left(\rho\right)H_{0}\left(\rho\right)\right]$